let – 1≤ p ≤1 .Show that the equation 4x3-3x-p=0 has a unique root in the interval [-1/2, 1] and identify it. 
Ans: x= cos {1/3 cos-1(p)}

Grade:Select Grade

2 Answers

Vikas TU
7 years ago
First find f(-0.5) and f(1) in terms of p.
And suit the domain for p lyng in [-1,1]
and check relation in  f(-0.5) and f(1).
Next step is to proove the function is continuous and differntiable.
I would be and then apply Mean Value theorem to it. simply.
Ajay
7 years ago
Firstly the question is wrong , the interval in which root will be unique is [1/2, 1] instead of [-1/2, 1].
The equation can be written as 4x3-3x  = p
Draw the graph of of function y = f(x) = 4x3-3x. by finding its maxima and minima.
The given equation will have unique root in any interval if line y = p cuts the graph y=f(x) at exactly one point.
From the graph it would be clear that that this will be true if interval of x  is [1/2, 1]
 
Let the root be a, so it will satisfy the given equation we have
4a3-3a = p
let a = cosr
4cos3r-3cosr = p
cos3r = p
r  = 1/3cos-1p
a = cosr = cos(1/3cos-1p) will be the required root